OpenStudy (anonymous):

a contractor builds two patios with the same area. one is a square and the other is a rectangle 10m longet than the square but 8m less wide than the square. what is the area of each patio?

OpenStudy (raden):

let the length of that square is x m, therefore the length and the width of rectangle respectively be (x+10) and (x-8). according information above the area of both is same, it means : Area of square = area of rectangle x * x = (x+10)(x-8) x^2 = x^2 + 2x - 80 2x - 80 = 0 2x = 80 so, x = 40

OpenStudy (raden):

after we got x, now we subtitute x=40 to the formula of each area of square and rectangle. area of square = x * x = 40 * 40 = 1600

OpenStudy (raden):

obviously, the area of rectangle = 1600 too because they are have the same area :)

OpenStudy (anonymous):

thank you so much can you do this one plz a rectangle is 12cm longer than it is wide. if its lenght and width are both decreased by 2cm. its area is decreased by 108 square cm. find its original dimensions show your work plz HELP!!!